### v1

@@
Abstract
@@
Explainable Artificial Intelligence (XAI) techniques, such as SHapley Additive exPlanations (SHAP), have become essential tools for interpreting complex ensemble tree-based models, especially in high-stakes domains such as healthcare analytics. However, SHAP values are usually treated as point estimates, which disregards the inherent and ubiquitous uncertainty in predictive models and data. This uncertainty has two primary sources: aleatoric and epistemic. The aleatoric uncertainty, which reflects the irreducible noise in the data. The epistemic uncertainty, which arises from a lack of data. In this work, we propose an approach for decomposing uncertainty in SHAP values into aleatoric, epistemic, and entanglement components. This approach integrates Dempster-Shafer evidence theory and hypothesis sampling via Dirichlet processes over tree ensembles. We validate the method across three real-world use cases with descriptive statistical analyses that provide insight into the nature of epistemic uncertainty embedded in SHAP explanations. The experimentations enable to provide more comprehensive understanding of the reliability and interpretability of SHAP-based attributions. This understanding can guide the development of robust decision-making processes and the refinement of models in high-stakes applications. Through our experiments with multiple datasets, we concluded that features with the highest SHAP values are not necessarily the most stable. This epistemic uncertainty can be reduced through better, more representative data and following appropriate or case-desired model development techniques. Tree-based models, especially bagging, facilitate the effective quantification of epistemic uncertainty.
keywords:
[…]
1 Introduction
@@
Machine learning (ML) [1] is a key part of improving healthcare analytics such as resource planing, disease diagnosis, prognosis, and risk stratification [2, 3, 4]. However powerful, uncertainty is inherent and ubiquitous in machine learning (ML) models because their predictions are affected by noisy data, model limitations, and unseen scenarios. To address this challenge, some of the most widely used tools are ensemble tree-based models [5], which help in managing and quantifying uncertainty in predictions. They are highly accurate, interpretable, and efficient with structured data, resulting in lower computational demand. Unlike deep neural networks, which require large amounts of unstructured data such as images and text, they have lower computational requirements and are more interpretable [6, 7]. These include Random Forest (RF) [8], Gradient Boosting Machines (GBM), and Extreme Gradient Boosting (XGBoost) [9]. These models are robust against noise, and can handle large, complicated data sets, which are common in healthcare [10]. Ensemble tree approaches are different from traditional machine learning models [11] as they can efficiently capture complex, nonlinear relationships and slight interactions among the features [12]. This leads to highly accurate and generalizable predictions. Ensemble models have many advantages. They combine the strengths of multiple base learners which reduces overfitting, improves stability, and enhances the model’s ability to generalize to unseen data unlike traditional ML models [7]. It has been shown that using a group of classifiers to make predictions outperforms using individual classifiers to predict heart disease [10, 13]. This is important for making reliable decisions, which makes them very useful in healthcare.Despite their strengths, ensemble tree-based models have two main problems. First, they are difficult to interpret, particularly when there are large numbers of constituent trees and features [14, 15, 16]. This “black box” nature, avoiding them to be the first choice in the model selection. To address this, explainable AI (XAI) techniques such as SHapley Additive exPlanations (SHAP) or Shapley values [17], which are rooted in cooperative game theory, have emerged as a principled framework for attributing the contribution of each feature to individual predictions in ML models. However, calculating Shapley values can be difficult for complex models. Fortunately, a couple of new, efficient methods for calculating them for certain types of models have recently emerged. TreeSHAP [18, 19], a fast and exact method for calculating Shapley values in tree-based models like decision trees, RFs, XGBoost, have been introduced. TreeSHAP uses the natural structure of decision trees to make predictions that are much faster and easier to understand than those of other methods. This is helpful for explaining results with large groups of data, which is important in fields where understanding predictions is paramount, for instance in healthcare or finance. SHAP values provide a clear framework for determining how each feature contributes to individual predictions. They offer insight into the decision-making process behind complex ensemble models. SHAP and similar methods not only encourage trust, but also make it easier to use advanced machine learning (ML) models in healthcare by making them easier to comprehend.
@@
Recent research has identified several factors that improve SHAP values. These factors include misattribution of feature importance, reliance on the assumption of feature independence, lack of causal or contextual understanding, computational inefficiency, and risk of misinterpretation. To address these issues, alternative attribution methods and error quantification techniques, such as Normalized Movement Rate (NMR) and Modified Index Position (MIP), have been proposed to handle feature collinearity [20]. New SHAP variants have also been developed, with the aim of improving efficiency and interpretability. Latent SHAP [21, 19] extends SHAP by enabling explanations in human-interpretable domains without requiring invertible transformations, making it suitable for high-dimensional or non-invertible data and capturing correlations among features. Kernel SHAP [22], the most versatile black-box SHAP explainer, uses weighted linear regression to approximate Shapley values and is valued for its generality. However, it is slower and assumes feature independence, which can limit accuracy in correlated data. Muschalik et al. [23] introduce methods for efficiently computing higher-order Shapley interactions in tree ensembles. This enables richer, more granular explanations of feature interactions than standard SHAP, with significant computational advantages for large or complex models. These advancements collectively enhance the reliability, interpretability, and practicality of SHAP-based explanations in ML. Other advancements include integrating causal and contextual information, as well as creating more computationally efficient SHAP variants, such as CF-SHAP and FF-SHAP [24].
@@
However, SHAP is a point-estimate method, which contributes to the uncertainty of the explanations it produces. This opens up new dimensions for studying and quantifying uncertainties in XAI, which is an important step forward in the field [25, 26, 27]. Current implementations, as discussed before, treat SHAP or TreeSHAP values with respect to tree-based ML models as point estimates. However, this approach ignores the individual contributions of epistemic uncertainty arising from variability in model training. It focuses on overall uncertainty, comprising aleatoric and epistemic uncertainty. Aleatoric uncertainty refers to the inherent randomness or noise in the data that cannot be reduced by collecting more information. Epistemic uncertainty, on the other hand, arises from a lack of knowledge or data about the model or process. This omission poses critical risks in high-stakes domains. For instance, medical diagnostics using XGBoost may yield identical SHAP values across hospitals despite shifts in regional data distribution. Similarly, financial risk models may exhibit unstable feature attributions during market volatility [28, 29]. Random forest (SHAP) models used for predictive health monitoring [30, 31] may appear similar or stable when applied to hospitals with different demographics or disease prevalence rates but SHAP values often exhibit bias toward features with higher cardinality or entropy. This bias can overstate or understate the importance of these features when patient populations change. Quantifying epistemic uncertainty in SHAP values is necessary because people trust model explanations for decision support, especially in high-stakes domains such as healthcare and finance [32, 33, 25]. Contemporary methods use techniques such as bootstrap sampling to estimate the uncertainty of SHAP attributions and generate confidence intervals for the importance of each feature. Now, variants of SHAP enable users to evaluate the reliability of feature contributions instead of relying exclusively on point estimates. These methods allow practitioners to more accurately evaluate the stability of explanations, identify features or contexts in which explanations are less reliable, and improve model transparency in situations involving shifting or uncertain data.
@@
    Decomposes SHAP variance into aleatoric and epistemic components
@@
    Leverages belief functions and Dirichlet processes for hypothesis space sampling
@@
2 Background
2.1 SHAP
@@
Shapley values enable the quantification and interpretation of feature contributions in ML. Research shows SHAP as one of the most interpretable methods in ML, providing insight into complex healthcare ML models. It is a model-agnostic interpretability tool used extensively in healthcare. [34, 35, 36, 37]. Furthermore, it has been used in fields such as predicting breast cancer risk, elucidating model predictions, diagnosing biomarkers, and analyzing survival, particularly with tree-based machine learning models [38, 39, 10]. SHAP has been used to interpret machine learning models that predict cancer risk. For example, it has identified age and family history as key predictors of breast cancer risk [38]. Furthermore, SHAP has improved the interpretability of models for other chronic diseases. It has been used to examine machine learning models for smoking and drinking habits, using lifestyle data, blood test results, and wearable sensor readings. This has facilitated the interpretation of key influencing features and enhanced transparency for potential use in personalized healthcare [40]. In use cases involving imaging or clinical data, SHAP reveals the importance of diagnostic features. For example, SHAP emphasizes the texture and morphology of tumors in breast cancer mammography and the key variables in detecting chronic diseases[41]. The field of radiology stands to benefit from SHAP because it facilitates the use of AI models to interpret imaging scans for abnormalities, such as lung nodules or diabetic retinopathy [42]. This development has the potential to improve both diagnostic accuracy and patient trust. In the context of retinoblastoma diagnosis, SHAP was employed to generate local and global interpretations, highlighting specific regions and features in fundus images that substantially influence the model’s predictions  [43]. SHAP was also used in deep learning models to identify image features that facilitate early cancer detection and enhance the interpretability of automated histopathology analyses [44]. Applying SHAP to deep learning (DL) models in medical image analysis provides clinicians with visual interpretations of model predictions. This improves the understanding and validation of automated diagnoses in various imaging tasks [45]. Recent frameworks continue to adopt SHAP as a technique for improving explainability in healthcare [46].
2.2 Uncertainty in XAI
@@
There is a growing interest and there have been recent advances which focus on developing methods for uncertainty quantification in XAI [47, 48]. These methods focus on communicating the uncertainty associated with the interpretations, which is necessary for the wider adoption of AI in high-stakes scenarios. The quantification of the uncertainty related to the interpretations involves the study of the change in interpretation when the input data or the model parameters are changed. One of the recent works introduces a framework that models the interpretations as a function $e_{\theta}(x,f)$ where for a model $f$, an instance $x$, and explanation parameters $\theta$, the explanation $e_{\theta}(x,f)$ quantifies each feature’s contribution to the prediction  [49]. The function allows researchers to follow the uncertainty from the inputs and the model with the help of interpretations. Methods like this one frequently use empirical and analytical estimations. The former often include Monte Carlo simulations techniques that enable researchers to obtain multiple different version of the input or the model, allowing them to study the variance of the model output and corresponding interpretations. The latter, focus on how small changes in the inputs or the model parameters affect the interpretations, allowing the creation of a co-variance matrix that quantifies the uncertainty. Another recent work [49] introduces the Mean Uncertainty in the Explanations (MUE) metric which summarizes the overall uncertainty by normalizing the trace of the interpretations’ co-variance matrix. The method also enables researchers to directly compare the uncertainty values across different methods and models. However, most of the studies mentioned earlier show that XAI methods are only point estimates. This means that they can identify important features, but they don’t show how reliable the interpretability results are when different inputs and model parameters are used [50].
@@
Some of the works have made significant achievements in the terms of quantifying the uncertainty of SHAP values by calculating the confidence intervals and distributions. This is crucial in decision-making in healthcare [51]. The way traditional SHAP scores are calculated is limited because they rely on input data that is either well-specified or estimated accurately [52]. The standard SHAP framework requires knowledge of the underlying probability distribution, which is often either unknown or estimated from a small number of samples. This can lead to unstable or misleading feature importance estimates. A recent framework [26], calculates the SHAP score as a function over a range of possible distributions. The approach provides tight intervals for feature importance and shows that feature rankings can be very sensitive to distributional assumptions. The framework also shows that related decision problems can’t be solved using computers. In other words, these problems are NP-complete. These problems include determining if a feature SHAP score can be higher than a threshold or always outperform another feature, even for decision trees. Studies have shown that SHAP intervals can be quite wide when there’s uncertainty about the data distribution [26]. However, as more data becomes available, these intervals become more stable.
@@
Researchers have come up with new ways to understand the predictions and the uncertainty of models. These new ways use information theory to explain the predictions and uncertainty. They also help reduce uncertainty in certain features and provide efficient algorithms and methods for making inferences in the real world. These research areas show how important it is to understand the assumptions about data distribution. This is important for making sure that the features used in explainable machine learning can be trusted and understood. In healthcare, incorporating uncertainty quantification has allowed researchers to develop models more reliable and transparent [53]. It is very important to combine uncertainty quantification with XAI in high-stakes application, especially when the people who know the most about the domain utilize both the model predictions and explanations to make decisions. Recent research [25, 54, 32] on quantifying uncertainty in SHAP has improved the interpretability of ML models. The research by Cohen et al. [32] and Watson et al. [25] are significant in terms of extending XAI methods based on SHAP values to better quantify and explain uncertainty in model predictions. Cohen et al. introduce efficient methods and visual tools for measuring uncertainty in stochastic SHAP explanations, making them more applicable to real-time or high-stakes scenarios. Watson et al. use information theory to expand SHAP values and quantify the predictive uncertainty, offering formal reliability guarantees and scalable algorithms for practical use. But, both the approaches are sensitive to data and model quality. They primarily address uncertainty from estimator sampling. The challenges in terms of assessing broader sources of uncertainty. Overall, these significant developments help make AI explanations more transparent and trustworthy, especially with regard to uncertainty. However, further research is needed to study uncertainty arising from models and data.
@@
Although these works enable users to evaluate feature importance and model attribution confidence, additional dimensions must be addressed. Current methods include uncertainty aggregation and assign non-zero importance to irrelevant features. They are also sensitive to data and model biases, computationally intensive, and may produce unreliable attributions under model instability. Additionally, they lead to misleading interpretations in high-stakes scenarios and cannot distinguish between aleatoric and epistemic uncertainty. These methods are also vulnerable to adversarial manipulation. In the context of healthcare, SHAP values may lead domain professionals to overconfidently rely on feature importance when making treatment decisions based on model explanations. They may ignore epistemic uncertainty because data can be limited, biased, or heterogeneous. Furthermore, SHAP does not represent aleatoric uncertainty due to noisy or ambiguous input data. Misleading explanations can also occur if the model is biased, the data is collinear, or the underlying relationships are not well captured. Decomposing SHAP uncertainty into epistemic and aleatoric components can address these limitations [55]. This allows for targeted model improvement and enhanced, actionable explanations. However, as commonly implemented, SHAP values do not account for epistemic uncertainty, which arises from variability in model training. This limitation arises from SHAP’s design as a post hoc explanation tool for individual predictions rather than as a method for quantifying uncertainty.
3 Problem Formulation
@@
Theorem 1.
For any tree ensemble model $f$, the point estimate SHAP value $\phi_{i}$ lacks a measure of variance V($\phi_{i}$|f, D) over possible training datasets $D\sim P_{\text{data}}$. This violates the reliability axiom for explainability in high-risk AI systems [56].
@@
3.1 SHAP Variance Decomposition
The conventional uncertainty quantification framework posits that aleatoric uncertainty stems from inherent data noise and that epistemic uncertainty stems from model ignorance. However, empirical evidence challenges this distinction by demonstrating that, under shifts in the data distribution or model misspecification, aleatoric and epistemic uncertainties become intertwined. Bootstrap ensembles and deep ensemble methods show that, as epistemic uncertainty increases, estimates of aleatoric uncertainty can decrease, leading to systematic bias in model predictions [57, 58, 59]. For any feature $i$ and instance $\mathbf{x}$, the total variance of SHAP values $\phi_{i}(\mathbf{x})$ over possible training datasets $D\sim P_{\text{data}}$ and tree ensemble models $f$ decomposes as:
$\underbrace{\text{Var}_{D,f}(\phi_{i})}_{\text{Total}}=\underbrace{\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|f)]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])}_{\text{Epistemic}}+\underbrace{2\cdot\mathcal{C}(f,D)}_{\text{Entanglement}}$ (1)
where $\mathcal{C}(f,D)=\text{Cov}_{P(f,D)}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)$, and $\phi_{i}(x)$ The terms in the equations can be defined as follows:
@@ 3.1 SHAP Variance Decomposition
We have assumed the following assumptions:
@@
    Finite Variance: $\text{Var}(\phi_{i}|f)$ and $\text{Var}(\mathbb{E}[\phi_{i}|D])$ exist $\forall f,D$

### v2

@@
UbiQTree: Uncertainty Quantification in XAI with Tree Ensembles
Akshat Dubey Center for Artificial Intelligence in Public Health Research (ZKI-PH), Robert Koch Institute, Nordufer 20, 13353 Berlin, Germany Department of Mathematics and Computer Science, Freie Universität Berlin, Arnimallee 14, 14195 Berlin, Germany Lead contact Aleksandar Anžel Bahar İlgen Georges Hattab Center for Artificial Intelligence in Public Health Research (ZKI-PH), Robert Koch Institute, Nordufer 20, 13353 Berlin, Germany Department of Mathematics and Computer Science, Freie Universität Berlin, Arnimallee 14, 14195 Berlin, Germany
SUMMARY
@@
Explainable artificial intelligence (XAI) techniques, particularly SHapley Additive exPlanations (SHAP), are essential for interpreting ensemble tree-based models in critical areas such as healthcare. However, SHAP values are often treated as point estimates that neglect uncertainty originating from aleatoric (irreducible noise) and epistemic (lack of data) sources. This work introduces an approach that decomposes SHAP value uncertainty into aleatoric, epistemic, and entanglement components. This approach employs Dempster–Shafer evidence theory and Dirichlet process hypothesis sampling over tree ensembles. The use-case validation reveals insights into epistemic uncertainty within SHAP explanations, enhancing the reliability and interpretability of SHAP attributions. This informs robust decision-making and model refinement. Our findings suggest that reducing epistemic uncertainty requires improved data quality and model development techniques. Tree-based models, particularly bagging, are effective in quantifying such uncertainties.
KEYWORDS
Machine Learning, Healthcare, XAI, Explainability, SHAP, Random Forest, Ensemble Machine Learning, Uncertainty, Statistics, Evidence Theory
INTRODUCTION
@@
Machine learning (ML) 34 is a key part of improving healthcare analytics such as resource planing, disease diagnosis, prognosis, and risk stratification 39, 44, 61. However powerful, uncertainty is inherent and ubiquitous in machine learning (ML) models because their predictions are affected by noisy data, model limitations, and unseen scenarios. To address this challenge, some of the most widely used tools are ensemble tree-based models 57, which help in managing and quantifying uncertainty in predictions. They are highly accurate, interpretable, and efficient with structured data, resulting in lower computational demand. Unlike deep neural networks, which require large amounts of unstructured data such as images and text, they have lower computational requirements and are more interpretable 68, 92. These include Random Forest (RF) 8, Gradient Boosting Machines (GBM), and Extreme Gradient Boosting (XGBoost) 11. These models are robust against noise, and can handle large, complicated data sets, which are common in healthcare 32. Ensemble tree approaches are different from traditional machine learning models 17 as they can efficiently capture complex, nonlinear relationships and slight interactions among the features 19. This leads to highly accurate and generalizable predictions. Ensemble models have many advantages. They combine the strengths of multiple base learners which reduces overfitting, improves stability, and enhances the model’s ability to generalize to unseen data unlike traditional ML models 92. It has been shown that using a group of classifiers to make predictions outperforms using individual classifiers to predict heart disease 32, 79. This is important for making reliable decisions, which makes them very useful in healthcare.Despite their strengths, ensemble tree-based models have two main problems. First, they are difficult to interpret, particularly when there are large numbers of constituent trees and features 43, 25, 23. This “black box” nature, avoiding them to be the first choice in the model selection. To address this, explainable AI (XAI) techniques such as SHapley Additive exPlanations (SHAP) or Shapley values 56, which are rooted in cooperative game theory, have emerged as a principled framework for attributing the contribution of each feature to individual predictions in ML models. The field of XAI is rapidly evolving. Few of the recent advancements includes Counterfactual Paths (CPATH) which identifies feature permutations that influence model predictions and uses domain knowledge graphs 69. Total Causal Effect Calculation for Fuzzy Cognitive Maps (TCEC-FCM) 36 uses graph traversal techniques to compute causal effects, improving system transparency. Reciprocal Human-Machine Learning (RHML) 67 fosters continuous learning between humans and AI models, improving model performance and decision-making. Other recent methods 7 includes Diverse Counterfactual Explanations (DiCE) 62 which provides counterfactual explanations, which align with human cognitive processes and are highly valued for understanding a model. Logic Tensor Networks (LTN) 5 which is a framework rooted in Neural-symbolic AI, designed for learning and logical reasoning. It facilitates interactive explainability and model revision when applied to XAI. Template System for Natural Language Explanations (TS4NLE) 18 is designed for the presentation and rendering of explanations derived from other XAI outputs, prioritizing human comprehensibility. It can process the structured output of any approach and generate Natural Language Explanations (NLE), which are then rendered into tailored natural language text using a template system. These frameworks aim to make AI systems more transparent and aligned with human understanding. Despite these advancements, SHAP remains a popular candidate for studying XAI due to its strong theoretical foundation and widespread adoption 7. However, calculating Shapley values can be difficult for complex models. Fortunately, a couple of new, efficient methods for calculating them for certain types of models have recently emerged. TreeSHAP 55, 37, a fast and exact method for calculating Shapley values in tree-based models like decision trees, RFs, XGBoost, have been introduced. TreeSHAP uses the natural structure of decision trees to make predictions that are much faster and easier to understand than those of other methods. This is helpful for explaining results with large groups of data, which is important in fields where understanding predictions is paramount, for instance in healthcare or finance. SHAP values provide a clear framework for determining how each feature contributes to individual predictions. They offer insight into the decision-making process behind complex ensemble models. SHAP and similar methods not only encourage trust, but also make it easier to use advanced machine learning (ML) models in healthcare by making them easier to comprehend.
@@
Recent research has identified several factors that improve SHAP values. These factors include misattribution of feature importance, reliance on the assumption of feature independence, lack of causal or contextual understanding, computational inefficiency, and risk of misinterpretation. To address these issues, alternative attribution methods and error quantification techniques, such as Normalized Movement Rate (NMR) and Modified Index Position (MIP), have been proposed to handle feature collinearity 47. New SHAP variants have also been developed, with the aim of improving efficiency and interpretability. Latent SHAP 58, 37 extends SHAP by enabling explanations in human-interpretable domains without requiring invertible transformations, making it suitable for high-dimensional or non-invertible data and capturing correlations among features. Kernel SHAP 10, the most versatile black-box SHAP explainer, uses weighted linear regression to approximate Shapley values and is valued for its generality. However, it is slower and assumes feature independence, which can limit accuracy in correlated data. Muschalik et al. 63 introduce methods for efficiently computing higher-order Shapley interactions in tree ensembles. This enables richer, more granular explanations of feature interactions than standard SHAP, with significant computational advantages for large or complex models. These advancements collectively enhance the reliability, interpretability, and practicality of SHAP-based explanations in ML. Other advancements include integrating causal and contextual information, as well as creating more computationally efficient SHAP variants, such as CF-SHAP and FF-SHAP 89.
@@
However, SHAP is a point-estimate method, which contributes to the uncertainty of the explanations it produces. This opens up new dimensions for studying and quantifying uncertainties in XAI, which is an important step forward in the field 88, 14, 81. Current implementations, as discussed before, treat SHAP or TreeSHAP values with respect to tree-based ML models as point estimates. However, this approach ignores the individual contributions of epistemic uncertainty arising from variability in model training. It focuses on overall uncertainty, comprising aleatoric and epistemic uncertainty. Aleatoric uncertainty refers to the inherent randomness or noise in the data that cannot be reduced by collecting more information. Epistemic uncertainty, on the other hand, arises from a lack of knowledge or data about the model or process. This omission poses critical risks in high-stakes domains. For instance, medical diagnostics using XGBoost may yield identical SHAP values across hospitals despite shifts in regional data distribution. Similarly, financial risk models may exhibit unstable feature attributions during market volatility 90, 26. Random forest (SHAP) models used for predictive health monitoring 54, 16 may appear similar or stable when applied to hospitals with different demographics or disease prevalence rates but SHAP values often exhibit bias toward features with higher cardinality or entropy. This bias can overstate or understate the importance of these features when patient populations change. Quantifying epistemic uncertainty in SHAP values is necessary because people trust model explanations for decision support, especially in high-stakes domains such as healthcare and finance 15, 48, 88. Contemporary methods use techniques such as bootstrap sampling to estimate the uncertainty of SHAP attributions and generate confidence intervals for the importance of each feature. Now, variants of SHAP enable users to evaluate the reliability of feature contributions instead of relying exclusively on point estimates. These methods allow practitioners to more accurately evaluate the stability of explanations, identify features or contexts in which explanations are less reliable, and improve model transparency in situations involving shifting or uncertain data.
@@
    Decomposes SHAP variance into aleatoric, epistemic, and entanglement components.
@@
    Leverages belief functions and Dirichlet processes for hypothesis space sampling.
@@
The experiments and complete code necessary for reproducibility are available at https://github.com/dubeyakshat07/UBiQTree 22
BACKGROUND
SHAP
@@
Shapley values enable the quantification and interpretation of feature contributions in ML. Research shows SHAP as one of the most interpretable methods in ML, providing insight into complex healthcare ML models. It is a model-agnostic interpretability tool used extensively in healthcare. 1, 70, 74, 82. Furthermore, it has been used in fields such as predicting breast cancer risk, elucidating model predictions, diagnosing biomarkers, and analyzing survival, particularly with tree-based machine learning models 33, 27, 32. SHAP has been used to interpret machine learning models that predict cancer risk. For example, it has identified age and family history as key predictors of breast cancer risk 33. Furthermore, SHAP has improved the interpretability of models for other chronic diseases. It has been used to examine machine learning models for smoking and drinking habits, using lifestyle data, blood test results, and wearable sensor readings. This has facilitated the interpretation of key influencing features and enhanced transparency for potential use in personalized healthcare 85. In use cases involving imaging or clinical data, SHAP reveals the importance of diagnostic features. For example, SHAP emphasizes the texture and morphology of tumors in breast cancer mammography and the key variables in detecting chronic diseases38. The field of radiology stands to benefit from SHAP because it facilitates the use of AI models to interpret imaging scans for abnormalities, such as lung nodules or diabetic retinopathy 24. This development has the potential to improve both diagnostic accuracy and patient trust. In the context of retinoblastoma diagnosis, SHAP was employed to generate local and global interpretations, highlighting specific regions and features in fundus images that substantially influence the model’s predictions  2. SHAP was also used in deep learning models to identify image features that facilitate early cancer detection and enhance the interpretability of automated histopathology analyses 87. Applying SHAP to deep learning (DL) models in medical image analysis provides clinicians with visual interpretations of model predictions. This improves the understanding and validation of automated diagnoses in various imaging tasks 80. Recent frameworks continue to adopt SHAP as a technique for improving explainability in healthcare 21.
Uncertainty in XAI
@@
There is a growing interest and there have been recent advances which focus on developing methods for uncertainty quantification in XAI 3, 28. These methods focus on communicating the uncertainty associated with the interpretations, which is necessary for the wider adoption of AI in high-stakes scenarios. The quantification of the uncertainty related to the interpretations involves the study of the change in interpretation when the input data or the model parameters are changed. One of the recent works introduces a framework that models the interpretations as a function $e_{\theta}(x,f)$ where for a model $f$, an instance $x$, and explanation parameters $\theta$, the explanation $e_{\theta}(x,f)$ quantifies each feature’s contribution to the prediction  12. The function allows researchers to follow the uncertainty from the inputs and the model with the help of interpretations. Methods like this one frequently use empirical and analytical estimations. The former often include Monte Carlo simulations techniques that enable researchers to obtain multiple different version of the input or the model, allowing them to study the variance of the model output and corresponding interpretations. The latter, focus on how small changes in the inputs or the model parameters affect the interpretations, allowing the creation of a co-variance matrix that quantifies the uncertainty. Another recent work 12 introduces the Mean Uncertainty in the Explanations (MUE) metric which summarizes the overall uncertainty by normalizing the trace of the interpretations’ co-variance matrix. The method also enables researchers to directly compare the uncertainty values across different methods and models. However, most of the studies mentioned earlier show that XAI methods are only point estimates. This means that they can identify important features, but they don’t show how reliable the interpretability results are when different inputs and model parameters are used 13.
@@
Some of the works have made significant achievements in the terms of quantifying the uncertainty of SHAP values by calculating the confidence intervals and distributions. This is crucial in decision-making in healthcare 75. The way traditional SHAP scores are calculated is limited because they rely on input data that is either well-specified or estimated accurately 4. The standard SHAP framework requires knowledge of the underlying probability distribution, which is often either unknown or estimated from a small number of samples. This can lead to unstable or misleading feature importance estimates. A recent framework 14, calculates the SHAP score as a function over a range of possible distributions. The approach provides tight intervals for feature importance and shows that feature rankings can be very sensitive to distributional assumptions. The framework also shows that related decision problems can’t be solved using computers. In other words, these problems are NP-complete. These problems include determining if a feature SHAP score can be higher than a threshold or always outperform another feature, even for decision trees. Studies have shown that SHAP intervals can be quite wide when there’s uncertainty about the data distribution 14. However, as more data becomes available, these intervals become more stable.
@@
Researchers have come up with new ways to understand the predictions and the uncertainty of models. These new ways use information theory to explain the predictions and uncertainty. They also help reduce uncertainty in certain features and provide efficient algorithms and methods for making inferences in the real world. These research areas show how important it is to understand the assumptions about data distribution. This is important for making sure that the features used in explainable machine learning can be trusted and understood. In healthcare, incorporating uncertainty quantification has allowed researchers to develop models more reliable and transparent 77. It is very important to combine uncertainty quantification with XAI in high-stakes application, especially when the people who know the most about the domain utilize both the model predictions and explanations to make decisions. Recent research 88, 46, 15 on quantifying uncertainty in SHAP has improved the interpretability of ML models. The research by Cohen et al. 15 and Watson et al. 88 are significant in terms of extending XAI methods based on SHAP values to better quantify and explain uncertainty in model predictions. Cohen et al. introduce efficient methods and visual tools for measuring uncertainty in stochastic SHAP explanations, making them more applicable to real-time or high-stakes scenarios. Watson et al. use information theory to expand SHAP values and quantify the predictive uncertainty, offering formal reliability guarantees and scalable algorithms for practical use. But, both the approaches are sensitive to data and model quality. They primarily address uncertainty from estimator sampling. The challenges in terms of assessing broader sources of uncertainty. Overall, these significant developments help make AI explanations more transparent and trustworthy, especially with regard to uncertainty. However, further research is needed to study uncertainty arising from models and data.
@@
Although these works enable users to evaluate feature importance and model attribution confidence, additional dimensions must be addressed. Current methods include uncertainty aggregation and assign non-zero importance to irrelevant features. They are also sensitive to data and model biases, computationally intensive, and may produce unreliable attributions under model instability. Additionally, they lead to misleading interpretations in high-stakes scenarios and cannot distinguish between aleatoric and epistemic uncertainty. These methods are also vulnerable to adversarial manipulation. In the context of healthcare, SHAP values may lead domain professionals to overconfidently rely on feature importance when making treatment decisions based on model explanations. They may ignore epistemic uncertainty because data can be limited, biased, or heterogeneous. Furthermore, SHAP does not represent aleatoric uncertainty due to noisy or ambiguous input data. Misleading explanations can also occur if the model is biased, the data is collinear, or the underlying relationships are not well captured. Decomposing SHAP uncertainty into epistemic and aleatoric components can address these limitations 45. This allows for targeted model improvement and enhanced, actionable explanations. However, as commonly implemented, SHAP values do not account for epistemic uncertainty, which arises from variability in model training. This limitation arises from SHAP’s design as a post hoc explanation tool for individual predictions rather than as a method for quantifying uncertainty.
PROBLEM FORMULATION
@@
Theorem 1
For any tree ensemble model $f$, the point estimate SHAP value $\phi_{i}$ lacks a measure of variance V($\phi_{i}$|f, D) over possible training datasets $D\sim P_{\text{data}}$. This violates the reliability axiom for explainability in high-risk AI systems 6.
@@
SHAP Variance Decomposition
The conventional uncertainty quantification framework posits that aleatoric uncertainty stems from inherent data noise and that epistemic uncertainty stems from model ignorance. However, empirical evidence challenges this distinction by demonstrating that, under shifts in the data distribution or model misspecification, aleatoric and epistemic uncertainties become intertwined. Bootstrap ensembles and deep ensemble methods show that, as epistemic uncertainty increases, estimates of aleatoric uncertainty can decrease, leading to systematic bias in model predictions 41, 35, 76. For any feature $i$ and instance $\mathbf{x}$, the total variance of SHAP values $\phi_{i}(\mathbf{x})$ over possible training datasets $D\sim P_{\text{data}}$ and tree ensemble models $f$ decomposes as:
$\underbrace{\text{Var}_{D,f}(\phi_{i})}_{\text{Total}}=\underbrace{\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])}_{\text{Epistemic}}-\underbrace{\mathcal{C}(f,D)}_{\text{Entanglement}}$ (1)
where the entanglement term is:
$\mathcal{C}(f,D)=\text{Cov}_{P(f,D)}\left(\mathbb{E}_{D}[\phi_{i}|f],\mathbb{E}_{f}[\phi_{i}|D]\right)$
The terms in the equations can be defined as follows:
@@
We assume the following:
@@
    Finite Moments: $\mathbb{E}_{D,f}[|\phi_{i}|^{2}]<\infty$, $\mathrm{Var}(\phi_{i}|f)$ and $\mathrm{Var}(\mathbb{E}[\phi_{i}|D])$ exist $\forall f,D$